OpenStudy (saiyan):

Algebra 2 Complex Numbers Help if you can Simplify the expression? -6+i/-5+i A. 31+i/26 B. 29+i/26 C.31+i/24 D.31+11i/26

Vocaloid (vocaloid):

multiply the numerator and denominator by the conjugate of the denominator

OpenStudy (saiyan):

What is the conjugate?

Vocaloid (vocaloid):

look at the denominator -5+i change the sign in the middle

OpenStudy (saiyan):

So is would be -5-i

Vocaloid (vocaloid):

right, the conjugate of -5+i is -5-i now, go back to the original expression (-6+i)/(-5+i) multiplying the numerator and denominator by (-5-i) gives us [(-5-i)(-6+i)] / [(-5+i)(-5-i)] simplify

Vocaloid (vocaloid):

hint: use the foil method on the numerator and denominator

OpenStudy (saiyan):

i got i^2+30+i/i^2+25

Vocaloid (vocaloid):

almost, check your denonimator again

OpenStudy (saiyan):

ok i checked it and it gave me the same thing for the denominator -i^2+25

Vocaloid (vocaloid):

right, there was a negative sign in front of i^2 that you missed before so we have (i^2+30+i)/(-i^2+25) we can rewrite this if we recognize that i^2 = -1 can you try substituting -1 for i^2 into your expression now?

OpenStudy (saiyan):

So i would replace -1 with all the i^2

Vocaloid (vocaloid):

yes, whenever you see "i^2" you can replace it with -1

OpenStudy (saiyan):

29+i/24

Vocaloid (vocaloid):

almost but not quite check your denominator again

Vocaloid (vocaloid):

-i^2+25 = ?

OpenStudy (saiyan):

did i mess up with the double negative in the denominator

Vocaloid (vocaloid):

yes

Vocaloid (vocaloid):

-i^2+25 = -(-1) + 25 = ?

OpenStudy (saiyan):

ok it would be B. 29+i/26

Vocaloid (vocaloid):

yeah! good job

OpenStudy (saiyan):

Thankyou for helping me

Vocaloid (vocaloid):

@Saiyan uh, small correction, I made a mistake with the numerator the numerator should be 30+i-i^2, giving is (31+i)/(26), sorry > >

OpenStudy (saiyan):

ok thx